# Robust $\mathcal{H}_{\infty}$ control and worst-case search in constrained parametric space

### v5

@@ II-B Convergence of sequential quadratic programming (SQP) for upper-$C^{1}$ functions
Proposition 3 extends the convergence result of SQP established in [28, Theorem 3.2] for differentiable objective functions to the case of upper-$C^{1}$ objective functions. Besides the upper-$C^{1}$ condition of $f$, the same assumptions as in [28, Theorem 3.2] are considered.
Proposition 3 (Convergence of SQP for upper-$C^{1}$ functions)
@@
(i) there exists $\alpha,\beta>0$ such that, for each $k$ and for any $x\in\mathbb{R}^{k}$:
@@
Proof.
Cf. Appendix. The proof follows that of [28, Theorem 3.2] and adapts it to the case where f is not differentiable but only upper-$C^{1}$. In particular, the upper semi-continuity of the Clarke subdifferential allows to maintain the argument linked to the continuity of the gradient in the original proof. The existence of an accumulation point of the solutions $p_{k}$ of the quadratic problems $Q(x_{k},H_{k},g_{x_{k}})$ also follows from the submonotonicity of the subdifferential of lower-$C^{1}$ functions [39] (cf. Lemma 1 in Appendix). ∎
@@
The purpose of our arguments for using smooth optimization is not to claim superiority over specialized nonsmooth optimization methods (for example [25, 24] for lower- or upper-$C^{1}$ functions specifically). Instead, we emphasize that readily available algorithms (e.g., in Matlab or Python toolboxes) can be applied successfully.
III Worst case search in constrained parametric space
III-A Problem formulation
In this section, the controller $K$ is fixed and, thus, voluntarily omitted in the notations.
As in [3], worst-case stability is studied by maximizing the spectral abscissa of the closed-loop matrix $A(\delta)$. The problem (4) can also be written as the following constrained optimization problem:
$\begin{array}[]{lll}&\underset{\delta\in\mathbb{R}^{k}}{\text{minimize}}&a(\delta):=-\alpha(A(\delta))\\ &\text{subject to }&-1\leq\delta_{i}\leq 1\\ &&c_{i}(\delta)\leq 0\end{array}$ (12)
Similarly for the performance problem (3), we define for the $\mathcal{H}_{\infty}$ norm of a generic transfer function $T_{zw}(s,\delta)$:
@@ III-A Problem formulation
and for the (squared) $\mathcal{H}_{2}$ norm:
[…]
III-B Properties of the objective functions
In this section, we consider a generic transfer function written as the Linear Fractional Transformation (LFT) [41, Chapter 10] $T_{zw}(s,\delta)=\mathcal{F}_{u}(M(s),\Delta)$, as depicted in Fig. 1, where $\Delta=\textrm{diag}(\delta_{i}I_{n_{i}})$ ($n_{i}$ is the number of repetitions of $\delta_{i}$). This representation will be used to characterize either the internal stability, or the $\mathcal{H}_{\infty}$ or $\mathcal{H}_{2}$ norm of the transfer from $w$ to $z$; practical problems e.g. (2) can also be addressed after an appropriate partitioning.
@@ III-B Properties of the objective functions
$\left\{\begin{array}[]{l}\left[\begin{array}[]{c}\dot{x}\\ \hline\cr q\\ z\end{array}\right]=\left[\begin{array}[]{c|cc}A_{1}&B_{1}&B_{2}\\ \hline\cr C_{1}&D_{11}&D_{12}\\ C_{2}&D_{21}&D_{22}\end{array}\right]\left[\begin{array}[]{c}x\\ \hline\cr p\\ w\end{array}\right]\\ p=\Delta q\end{array}\right.\iff\left\{\begin{array}[]{l}\left[\begin{array}[]{c}\dot{x}\\ \hline\cr\tilde{q}\\ z\end{array}\right]=\left[\begin{array}[]{c|cc}A_{1}&B_{1}&B_{2}\\ \hline\cr C_{1}&0&D_{12}\\ C_{2}&D_{21}&D_{22}\end{array}\right]\left[\begin{array}[]{c}x\\ \hline\cr\tilde{p}\\ w\end{array}\right]\\[15.0pt] \tilde{p}=\widetilde{\Delta}\tilde{q}\end{array}\right$.
where, in the right equation, the system was rewritten with the change of variable:
$\widetilde{\Delta}=(I-\Delta D_{11})^{-1}\Delta=\Delta(I-D_{11}\Delta)^{-1}$ (15)
to have no direct transfer from $\tilde{p}$ to $\tilde{q}$. This change of variable is the same as in [41, Section 12.3.4], with the uncertainty block instead of a feedback controller, and will be useful to study the $\mathcal{H}_{2}$ norm.
III-B1 $\mathcal{H}_{2}$ norm (function $h_{2}$)
Proposition 4 (Differentiability of $h_{2}$)
Let us assume $D_{22}=0$, and either $D_{12}=0$ or $D_{21}=0$. Then, $h_{2}$ is defined and differentiable in
$\mathcal{D}_{s}=\{\delta\in\mathcal{D},T_{zw}(s,\delta)\text{ is internally stable}\}$
and its gradient is given in Eq. (18).
While the proof mainly follows previously established results cited thereafter, it is necessary to present the following equations to obtain the explicit gradient expression (18). The assumption $D_{22}=0$ is necessary to ensure finiteness of the $\mathcal{H}_{2}$ norm in the nominal case $\Delta=0$. The assumption $D_{12}=0$ or $D_{21}=0$ is done for the same reason for $\Delta\neq 0$. The differentiability of $h_{2}$ in $\mathcal{D}_{s}=\{\delta\in\mathcal{D},T_{zw}(s,\delta)\text{ is stable}\}$ was shown in [34, Section 3]; it was originally derived for a controller rather than an uncertainty block, but the validity of the result extends straightforwardly to the latter. Let us write the state-space matrices of the transfer from $w$ to $z$ after closing the loop between $p$ and $q$:
$A(\delta)=A_{1}+B_{1}\widetilde{\Delta}C_{1}\;,\;B(\delta)=B_{2}+B_{1}\widetilde{\Delta}D_{12}\;,\;C(\delta)=C_{2}+D_{21}\widetilde{\Delta}C_{1}\;,\;D(\delta)=D_{21}\widetilde{\Delta}D_{12}=0$ (16)
Then, the squared $\mathcal{H}_{2}$ norm is:
$T_{zw}(s,\delta)||_{2}^{2}=\text{Tr}\left(B(\delta)^{T}X(\delta)B(\delta)\right)=\text{Tr}\left(C(\delta)Y(\delta)C(\delta)^{T}\right)$
where $X(\delta)$ and $Y(\delta)$ are solutions of the Lyapunov equations:
$A(\delta)^{T}X(\delta)+X(\delta)A(\delta)+C(\delta)^{T}C(\delta)$ $=0$
$A(\delta)Y(\delta)+Y(\delta)A(\delta)^{T}+B(\delta)B(\delta)^{T}$ $=0\;.$
The derivative of $h_{2}$ at point $\delta$ for the variable $\widetilde{\Delta}$, noted $h_{2}^{\prime}(\widetilde{\Delta})d\widetilde{\Delta}$, is obtained from [6, Lemma 3.1] or equivalently [12, Section 2.1], also originally derived for a controller and adapted here for the uncertain block $\widetilde{\Delta}$:
[…]
where the gradient $\nabla_{\widetilde{\Delta}}h_{2}(\delta)$ is given in the same references:
$\nabla_{\widetilde{\Delta}}h_{2}(\delta)=-2\left(B_{1}^{T}X(\delta)+D_{21}^{T}C(\delta)\right)Y(\delta)C_{1}^{T}-2B_{1}^{T}X(\delta)B(\delta)D_{12}^{T}\;.$
Finally, differentiating Eq. (15):
$d\widetilde{\Delta}=(I-\Delta D_{11})^{-1}d\Delta\left(I+D_{11}(I-\Delta D_{11})^{-1}\Delta\right)$
and using the same notation as in [3]:
$\Delta_{i}:=\frac{\partial\Delta}{\partial\delta_{i}}\;,$ (17)
the derivatives of $h_{2}$ read:
$\frac{\partial h_{2}}{\partial\delta_{i}}=-2\text{Tr}\left(\nabla_{\widetilde{\Delta}}h_{2}(\delta)^{T}(I-\Delta D_{11})^{-1}\Delta_{i}\left(I+D_{11}(I-\Delta D_{11})^{-1}\Delta\right)\right)\;.$ (18)
III-B2 $\mathcal{H}_{\infty}$ norm (function $h_{\infty}$)
Proposition 5 (Regularity of $h_{\infty}$ [16, 8, 3])
@@ III-B2 $\mathcal{H}_{\infty}$ norm (function $h_{\infty}$)
From the results of Section II and Proposition 5.b), we deduce that we can use SQP (Algorithm 1) to solve Problem (13). In particular, let us define the direction $d_{\delta}$ such that its components are defined as in Eq. (19) (even when $h_{\infty}$ is not differentiable), i.e. the i-th component of $d_{\delta}$ reads:
@@
where $\omega_{0}$ is an active frequency – but not necessarily the only one, and the multiplicity of $\bar{\sigma}(T_{zw}(j\omega_{0},\delta))$ is not necessarily 1 – and $p_{\omega_{0}}$ and $q_{\omega_{0}}$ are the corresponding normalized right and left singular vectors of $T_{zw}(j\omega_{0},\delta)$. Then, $d_{\delta}$ is a subgradient of $h_{\infty}$ (this is straightforward from the expression of the subdifferential given in [3, Section IV-1]); by Propositions 2 and 3, it can legitimately be used in SQP as if it were the gradient of a differentiable function.
@@
In contrast, minimizing the lower-$C^{1}$ function $h_{\infty}(\delta)$, i.e. minimizing the $\mathcal{H}_{\infty}$ norm, would be a more complex, genuinely nonsmooth problem. In particular, an unconstrained minimum may occur at a point where many subgradients are not zero (as a more simple example, consider $f:x\rightarrow|x|$, which is also lower-$C^{1}$). Moreover, descent directions are not provided by any subgradient. For this problem, nonsmooth techniques in the literature include multidirectionnal search [7], gradient sampling [18], or Clarke subdifferential-based [8, 15]. This is also pointed out in [3] as a fundamental difference between min-max and min-min (or max-max) programs, the former (e.g. minimizing the $\mathcal{H}_{\infty}$ norm) being genuinely non-smooth while the latter (e.g. maximizing the $\mathcal{H}_{\infty}$ norm or equivalently minimizing $h_{\infty}$) having the favorable properties demonstrated above.
III-B3 Spectral abscissa (function $a$)
@@
c) If $A(\delta)$ has only one active eigenvalue and it is simple, then $a$ is differentiable at $\delta$. With the notation of equations (16) and (17), the derivative of $a$ with regard to $\delta_{i}$ reads:
$\frac{\partial a}{\partial\delta_{i}}=-\mathrm{Tr}\left(\Delta_{i}^{T}\mathrm{Re}((I-D_{11}\Delta)^{-1}C_{1}v_{i}u_{i}^{H}B_{1}(I-\Delta D_{11})^{-1})^{T}\right)$ (20)
where $v_{i}$ and $u_{i}$ are right and left eigenvectors associated with the active eigenvalue, normalized such that $u_{i}^{H}v_{i}=1$.
Proposition 6.a) was established in [3] following the earlier work of [14]. These references build on the study of [19] concerning the directional derivatives of the spectral abscissa. b) is also shown in [3], and simply follows from the fact that the spectral abscissa is the maximum of the real parts of the eigenvalues, which are differentiable when the eigenvalues are simple [32]. For c), the expression (20) is straightforward from the expression of the Clarke subdifferential given in [3].
@@
In this paper, as for $h_{\infty}$, we rely on the upper-$C^{1}$ property of $a$ of Proposition 6.b) and on the results of Section II, when active eigenvalues are simple. However, the case of non-simple active eigenvalues is admittedly a limitation of using smooth optimization algorithms for the worst-case search in stability.
III-C Global exploration, local exploitation
Problems (12), (13) and (14) are constrained optimization problems involving the nonsmooth functions $a$ and $h_{\infty}$, and the differentiable function $h_{2}$. As demonstrated in Section II, Sequential Quadratic Programming may be used, as it will converge to KKT points of the optimization problems under the upper-$C^{1}$ assumption. This constitutes the main theoretical contribution of the present paper. From a practical perspective, however, we additionally investigate the use of Monte-Carlo sampling to perform a global exploration of the parameter space, with SQP subsequently ensuring efficient local exploitation.
III-C1 Sequential Quadratic Programming (SQP)
SPQ is described in Algorithm 1 and, as discussed in Section II, may be used to minimize upper-$C^{1}$ functions. In our numerical application, we used Matlab routine fmincon [22]. In fmincon, the matrix $H$ of the quadratic problem (10) serves as an approximation of the Hessian matrix of the Lagrangian, iteratively updated with the Broyden-Fletcher-Goldfarb-Shanno (BFGS) method. It is worth noting that BFGS was observed to perform well in [14] even for a non-smooth objective function (the spectral abscissa).
III-C2 Monte-Carlo (MC) sampling
Monte Carlo sampling is applied as follows: randomly pick a sample $\delta$ in the unit hypercube; evaluate $c(\delta)$: if positive, discard $\delta$, otherwise, evaluate the function $a(\delta)$ (respectively $h_{\infty}(\delta)$ or $h_{2}(\delta)$); repeat until having $N$ function evaluations. Since this paper addresses deterministic worst-case search, it is natural to draw parameters from uniform distributions; note, however, that removing some inadmissible configurations implicitly modifies these distributions. To leverage the local search enabled by SQP, we envision three ways to perform the Monte-Carlo sampling.
First, assume a probability $p$ that a random sample lies in the basin of attraction of the global optimum, i.e. that a gradient-based local optimization starting from this sample converges to the global optimum. If $N$ samples are drawn and all serve as initial point for an SQP run, the probability of finding the global optimum is $1-(1-p)^{N}$. Thus, given a target probability $p$ (reflecting the size of the basin of attraction of the global optimum that we wish to be able to detect), and a confidence level $P$ such that $1-(1-p)^{N}\geq P$, the required number $N$ of samples is

### v6

@@ 2.2 Sequential quadratic programming (SQP) algorithm for upper-$C^{1}$ functions
To prove the convergence of Algorithm 1, the following lemmas are necessary.
Lemma 5.
: Supermonotonicity of the subdifferential of upper-$C^{1}$ functions
Let $f$ be upper-$C^{1}$ at $x$. For every $\epsilon>0$, there exists $\delta>0$ such that for all $x_{0},x^{\prime}_{0}\in\mathcal{B}(x,\delta)$:
[…]
where $g_{x_{0}}\in\partial f(x_{0})$ and $g_{x^{\prime}_{0}}\in\partial f(x^{\prime}_{0})$.
{pf}
This comes from [37, Theorem 3.9] (submonotonicity of the subdifferential of lower-$C^{1}$ functions) and is also noted for example in [5, Lemma 6].
Lemma 6.
: Perturbation of a quadratic problem
Let us consider the two quadratic problems:
$x_{0}=\arg\min\{\;Q(x)=\frac{1}{2}x^{T}Kx-x^{T}k\;\;\\ \text{subject to }\;Gx\leq g\}$
$x_{0}^{\prime}=\arg\min\{\;Q^{\prime}(x)=\frac{1}{2}x^{T}K^{\prime}x-x^{T}k^{\prime}\;\;\\ \text{subject to }\;G^{\prime}x\leq g^{\prime}\}$
[…]
$x_{0}^{\prime}-x_{0}||\leq c\epsilon$
whenever $\epsilon<\epsilon_{0}$.
{pf}
Cf. Appendix. The proof is a mere adaptation of [22].
Proposition 7 extends the convergence result of SQP established in [27, Theorem 3.2] for differentiable objective functions to the case of upper-$C^{1}$ objective functions, when providing any subgradient where a gradient would be expected. Besides the upper-$C^{1}$ condition of $f$, the same assumptions as in [27, Theorem 3.2] are considered.
Proposition 7.
: Convergence of SQP for upper-$C^{1}$ functions
@@
(i) there exists $\alpha,\beta>0$ such that, for each $k$ and for any $x\in\mathbb{R}^{n}$:
@@
{pf}
The complete proof is provided in Appendix, and adapts that of [27, Theorem 3.2] to the case where f is not differentiable but only upper-$C^{1}$. Here, let us simply summarize said adaptations. In the proof of [27]:
1) it is always possible to choose $x_{k+1}$ as described in Algorithm 1 thanks to [27, Theorem 3.1]. This argument is maintained thanks to Proposition 3.a), which extends [27, Theorem 3.1] from differentiable to upper-$C^{1}$ functions.
2) given $\bar{x}$ be an accumulation point of $(x_{k})$, the continuity of the gradient of $f$ ensures that $\nabla f(x_{k})$ converges to $\nabla f(\bar{x})$. Here, a similar property is maintained as follows. Since $f$ is locally Lipschitz, its subgradients $(g_{x_{k}})$ are uniformly bounded for $k$ sufficiently large, hence there exists a subsequence that converges to a vector $\bar{g}$. Without loss of generality (passing to the subsequence if necessary) we note
$g_{x_{k}}\rightarrow\bar{g}\;,$
and the upper semi-continuity of the Clarke subdifferential [19, Proposition 1.5] yields $\bar{g}\in\partial f(\bar{x})$.
3) given the convergence of the (sub)sequences ($x_{k}$), ($H_{k}$), ($g_{x_{k}}$) and ($\nabla c_{i}(x_{k})$), the existence of an accumulation point of the solutions $p_{k}$ of the quadratic problems $Q(x_{k},H_{k},g_{x_{k}})$ follows from the continuity of $\nabla f(x)$ and from [22]. This argument is maintained in the upper-$C^{1}$ case thanks to Lemmas 5 and 6, where $x_{k+1}$ and $x_{k}$ play the roles of $x_{0}$ and $x^{\prime}_{0}$, and $g_{x_{0}}$ and $g_{x^{\prime}_{0}}$ play the roles of $k$ and $k^{\prime}$.
@@
The purpose of our arguments for using smooth optimization is not to claim superiority over specialized nonsmooth constrained optimization methods (for example [23, 24] for lower- or upper-$C^{1}$ functions specifically). Instead, we emphasize that readily available algorithms (e.g., in Matlab or Python toolboxes) can be applied successfully.
3 Worst-case search in constrained parametric space
3.1 Problem formulation
In this section, the controller $K$ is fixed and, thus, voluntarily omitted in the notations. For the performance problem (3), we define for the $\mathcal{H}_{\infty}$ norm of a generic transfer function $T_{zw}(s,\delta)$:
@@ 3.1 Problem formulation
We do not address the $\mathcal{H}_{2}$ norm in this section, because it is differentiable wherever it is well defined (cf. [32, 6, 11]) and therefore does not introduce any particular difficulty. As in [3], worst-case stability is studied by maximizing the spectral abscissa of the closed-loop matrix $A(\delta)$. Problem (4) can also be written as the following constrained optimization problem:
$\begin{array}[]{lll}&\underset{\delta\in\mathbb{R}^{n}}{\text{minimize}}&a(\delta):=-\alpha(A(\delta))\\ &\text{subject to }&-1\leq\delta_{i}\leq 1\\ &&c_{i}(\delta)\leq 0\end{array}$ (13)
3.2 Properties of the objective functions
In this section, we consider a generic transfer function written as the Linear Fractional Transformation (LFT) [40, Chapter 10] $T_{zw}(s,\delta)=\mathcal{F}_{u}(M(s),\Delta)$, as depicted in Fig. 1, where $\Delta=\textrm{diag}(\delta_{i}I_{n_{i}})$ ($n_{i}$ is the number of repetitions of $\delta_{i}$). This representation will be used to characterize either the internal stability or the $\mathcal{H}_{\infty}$ norm of the transfer from $w$ to $z$.
@@ 3.2 Properties of the objective functions
$\begin{array}[]{l}\left[\begin{array}[]{c}\dot{x}\\ \hline\cr q\\ z\end{array}\right]=\left[\begin{array}[]{c|cc}A_{1}&B_{1}&B_{2}\\ \hline\cr C_{1}&D_{11}&D_{12}\\ C_{2}&D_{21}&D_{22}\end{array}\right]\left[\begin{array}[]{c}x\\ \hline\cr p\\ w\end{array}\right]\end{array}$
Proposition 9.
: Regularity of $h_{\infty}$ [15, 8, 3]
@@
From the results of Section 2 and Proposition 9.b), we deduce that we can use SQP (Algorithm 1) to solve Problem (12). In particular, let us define the direction $d_{\delta}$ such that its components are defined as in Eq. (14) (even when $h_{\infty}$ is not differentiable), i.e. the i-th component of $d_{\delta}$ reads:
@@
where $\omega_{0}$ is an active frequency – but not necessarily the only one, and the multiplicity of $\bar{\sigma}(T_{zw}(j\omega_{0},\delta))$ is not necessarily 1 – and $p_{\omega_{0}}$ and $q_{\omega_{0}}$ are the corresponding normalized right and left singular vectors of $T_{zw}(j\omega_{0},\delta)$. Then, $d_{\delta}$ is a subgradient of $h_{\infty}$ (this is straightforward from the expression of the subdifferential given in [3, Section IV-1]); by Propositions 3 and 7, it can legitimately be used in SQP as if it were the gradient of a differentiable function.
@@
In contrast, minimizing the lower-$C^{1}$ function $h_{\infty}(\delta)$ (e.g. finding a controller that minimizes the $\mathcal{H}_{\infty}$ norm) would be a more complex, genuinely nonsmooth problem. In particular, an unconstrained minimum may occur at a point where many subgradients are not zero (as a more simple example, consider $f:x\rightarrow|x|$, which is also lower-$C^{1}$). Moreover, a descent direction cannot be found by picking any subgradient. For this problem, nonsmooth techniques in the literature include multidirectionnal search [7], gradient sampling [17], or Clarke subdifferential-based [8, 14]. This is also pointed out in [3] as a fundamental difference between min-max and min-min (or max-max) programs, the former (e.g. minimizing the $\mathcal{H}_{\infty}$ norm) being genuinely non-smooth while the latter (e.g. maximizing the $\mathcal{H}_{\infty}$ norm or equivalently minimizing $h_{\infty}$) having the favorable properties demonstrated above.
@@
c) If $A(\delta)$ has only one active eigenvalue and it is simple, then $a$ is differentiable at $\delta$. Noting $\Delta_{i}:=\frac{\partial\Delta}{\partial\delta_{i}}$, the derivative of $a$ with regard to $\delta_{i}$ reads:
$\frac{\partial a}{\partial\delta_{i}}=-\mathrm{Tr}(\Delta_{i}^{T}\mathrm{Re}((I-D_{11}\Delta)^{-1}C_{1}\\ v_{i}u_{i}^{H}B_{1}(I-\Delta D_{11})^{-1})^{T})$ (15)
where $v_{i}$ and $u_{i}$ are right and left eigenvectors associated with the active eigenvalue, normalized with $u_{i}^{H}v_{i}=1$.
Proposition 12.a) was established in [3] following the earlier work of [13]. These references build on [18] that studied the directional derivatives of the spectral abscissa. b) is also shown in [3], and simply follows from the fact that the spectral abscissa is the maximum of the real parts of the eigenvalues, which are differentiable when the eigenvalues are simple [31]. For c), the expression (15) is straightforward from the expression of the Clarke subdifferential given in [3].
@@
In this paper, as for $h_{\infty}$, we rely on the upper-$C^{1}$ property of $a$ of Proposition 12.b) and on the results of Section 2, when active eigenvalues are simple. However, the case of non-simple active eigenvalues is admittedly a limitation of using a smooth optimization algorithm for the worst-case search in stability. Nevertheless, this limitation is partially circumvented by the Monte-Carlo exploration proposed thereafter.
3.3 Global exploration, local exploitation
Problems (12) and (13) are constrained optimization problems involving the nonsmooth functions $h_{\infty}$ and $a$, which are upper-$C^{1}$ under the assumptions of Propositions 9 and 12. As demonstrated in Section 2, Sequential Quadratic Programming may be used, as it will converge to KKT points of the optimization problems. This constitutes the main theoretical contribution of this paper. From a practical perspective, however, we additionally investigate the use of Monte-Carlo sampling to perform a global exploration of the parameter space, with SQP subsequently ensuring efficient local exploitation. The advantages of this approach will be illustrated in the application of Section 5.
Local exploitation with Sequential Quadratic Programming (SQP)
SPQ is described in Algorithm 1 and, as discussed in Section 2, may be used to minimize upper-$C^{1}$ functions. A common choice for the matrix $H$ of the quadratic problem (10) is the Broyden-Fletcher-Goldfarb-Shanno (BFGS) method, which iteratively updates it to serve as an approximation of the Hessian matrix of the Lagrangian. It is worth noting that BFGS was observed to perform well in [13] even for a non-smooth objective function (the spectral abscissa).
Global exploration with Monte-Carlo sampling
Monte-Carlo sampling is applied as follows: randomly pick a sample $\delta$ in the unit hypercube; evaluate $c(\delta)$: if positive, discard $\delta$, otherwise, evaluate the function $a(\delta)$ (respectively $h_{\infty}(\delta)$ or $h_{2}(\delta)$); repeat until having $N$ function evaluations. Since this paper addresses deterministic worst-case search, it is natural to draw parameters from uniform distributions; note, however, that removing some inadmissible configurations implicitly modifies these distributions. To leverage the local search enabled by SQP, we envision different ways to perform the Monte-Carlo sampling.
First, assume a probability $p$ that a random sample lies in the basin of attraction of the global optimum, i.e. that a gradient-based local optimization starting from this sample converges to the global optimum. If $N$ samples are drawn and all serve as initial point for an SQP run, there is a probability $1-(1-p)^{N}$ of finding the global optimum. Thus, given a target probability $p$ (reflecting the size of the basin of attraction of the global optimum that we wish to be able to detect), and a confidence level $P$ such that $1-(1-p)^{N}\geq P$, the required number $N$ of samples is